OpenStudy (anonymous):

if x=5root2 - root6 then 56-20root3/5root2+root6 in terms of x is

OpenStudy (unklerhaukus):

\[x=5\sqrt2-\sqrt6\] \[56-\frac{20\sqrt3} {5\sqrt2}+\sqrt6\] ??

OpenStudy (anonymous):

no that root 6 is with 5root2

OpenStudy (anonymous):

\[(56-20\sqrt{3})x/44\]

OpenStudy (anonymous):

any one have doubts in question

OpenStudy (unklerhaukus):

yeah i still dont get the question

OpenStudy (anonymous):

i gave you the answer

OpenStudy (anonymous):

if x=52√−6√ 56−203√52√+6√ this we have to write in terms of x root6 is with 5root2

OpenStudy (anonymous):

options x^3/44 x/44

OpenStudy (anonymous):

x^3/44

OpenStudy (anonymous):

becouse (56−20 √ 3)=x^2

OpenStudy (anonymous):

x^3 ???

OpenStudy (anonymous):

ok can u plzzzz give the steps to do it plzzzz

OpenStudy (anonymous):

look my previous answer and add this comment and you get your answer

OpenStudy (anonymous):

steps are: multiply the denominator by conjugate, which is x and rest is just operations

OpenStudy (anonymous):

let me try it ......

OpenStudy (anonymous):

numerator how to solve it plzzz

OpenStudy (anonymous):

\[(56-20\sqrt{3})x/(50-6)=x ^{2}*x/44=x^{3}/44\]

OpenStudy (anonymous):

x*x=(5root2 - root6)(5root2 - root6)=56-20sqrt3

OpenStudy (anonymous):

how (56−203√)x become x^3

OpenStudy (anonymous):

x*x=(5root2 - root6)(5root2 - root6)=56-20sqrt3

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

hey the numerator i got as 300root2 - 156root6

OpenStudy (anonymous):

hmm... (5root2 - root6)(5root2 - root6)=25*2-5sqrt12-5sqrt12+6=56-10sqrt12=56-20sqrt3

OpenStudy (anonymous):

denominator is 44 i got it i want to knw about the numerator

OpenStudy (anonymous):

step1 multiply by conjugate: (56-20root3)(5root2-root6)/(5root2+root6)(5root2-root6) step 2: notice that: (5root2 - root6)(5root2 - root6)= 56-20sqrt3 step 3 substitute the step 2 expression in stpe 1 expretion: (5root2 - root6) (5root2 - root6)5root2-root6)/(5root2+root6)(5root2-root6) step 4 perform operation in denominator: (5root2 - root6) (5root2 - root6)5root2-root6)/44 step 5 substitute (5root2 - root6) =x in step 4 expretion: x^3/44

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

wait..

OpenStudy (anonymous):

okkk thanzxxxx

OpenStudy (anonymous):

the largest number that leaves reminder 7 abd 8 when divided by 3248 and 4175 respectively is??

OpenStudy (anonymous):

can u solve this

OpenStudy (anonymous):

sry got to go

OpenStudy (anonymous):

ohh